OpenStudy (anonymous):

Find the probability. A die is rolled 20 times and the number of twos that come up is tallied. Find the probability of getting exactly five twos. A. 0.129 B. 0.921 C. 0.003 D. 0.083

OpenStudy (anonymous):

binomial prob for this one \[P(X=k)=\dbinom{n}{k}p^k(1-p)^{n-k}\] in your example \(n=20,k=5,p=\frac{1}{6},1-p=\frac{5}{6}\)

OpenStudy (anonymous):

your answer is therefore \[\dbinom{20}{5}(\frac{1}{6})^5(\frac{5}{6})^{15}\]

OpenStudy (anonymous):

wats the answer

OpenStudy (anonymous):

i sent it to you